OpenStudy (swag):

how o u solve (3x^2y)^3

OpenStudy (mertsj):

\[(3x^2y)^3=3^3(x^2)^3(y^1)^3\]

OpenStudy (mertsj):

Does that help?

OpenStudy (mertsj):

What is 3 to the third?

OpenStudy (campbell_st):

the power of 3 is operating on each part of the expression and can be written in expanded form as \[3^3 \times (x^2)^3 \times (y^1)^3\] to simplify the variables use the index law for power of a power which means multiply the powers \[(x^a)^b = x^{a \times b}\]

OpenStudy (swag):

thank u guys